please help
Br Br 1) Excess NaNH2 ? 2) EtC 3) H2, Lindlars catalyst Draw the expected final product in the box below. Edit
1) NaNH2 2) Mel H CEC H ? 3) 9-BBN 4) H202, NaOH Draw the expected final product in the box below. Edit
1) NaNH2 2) Mel H CE C H 3) NaNH2 4) Et 5) Na, NH3() Draw the expected final product in the box below. ? Edit
1) NaNH2 2) Et H C C H ? 3) HgSO4 H2SO4. H2O Draw the expected final product in the box below. Edit
Answer #1

Answer By Br ① Excers NGNHz Et ② H2, Linollaus í catalyst E+ cis product f product (draw quy one of these CM2-CM3- H-C=come I 9 - BBN H-C=c-me BBN H-C=C-H 0 Nantz H-C=CH-CI I 202, Naom (4- 8 - CH2 - CH3 produer 3 H-C=C-H ① NGNM2H-C=comme